### v1

# Gaia Barium Dwarfs and Their Ostensibly Ordinary Counterparts

Abstract
The recently identified Gaia population of main-sequence–white dwarf (MS+WD) binaries at separations of , often with moderate eccentricities, is not readily reproduced by binary population synthesis models. Barium stars represent a closely related population whose enrichment in s-process elements both confirms the presence of a WD companion and attests to past binary interaction. It also indicates that mass transfer occurred at least during the late and post-AGB phases of the WD progenitor, when s-process elements are dredged up. In this work, we further explore the connection between the astrometrically identified Gaia MS+WD binaries and the classical barium star population. To achieve this, we used high-resolution FEROS spectroscopy to measure abundances for 30 Gaia DR3 non-single-star binaries, identifying 10 as Ba-enriched. Together with our recent analysis of archival GALAH data, this yields a sample of 38 barium dwarfs with dynamically measured WD masses, compared to only 6 previously known systems with known WD masses at these separations. We find that, in cases where metallicity is sufficiently low to facilitate efficient s-process production, barium and yttrium enrichment is often detected. This enrichment is also identified in eccentric systems, suggesting that post-AGB mass transfer mechanisms are capable of pumping eccentricity into the orbit or occur without erasing it. Our results indicate that the Gaia MS+WD binaries trace the population from which barium stars emerge. Treating the large Gaia-discovered population as an extension of known s-process enriched dwarfs opens an avenue to empirically constrain their formation and evolution.
1 Introduction
Barium stars (W. P. Bidelman & P. C. Keenan, 1951) were first discovered as red-giant stars of spectral type G-K with enhanced abundances of elements produced by the slow neutron capture process (s-process). Barium dwarfs are main-sequence (MS) stars of classes F-K with similar s-process element enhancement (P. North et al., 1994). This type of chemical enrichment has long been interpreted as a fossil record of past binary interactions.
Since s-process elements cannot be produced internally by stars before the asymptotic giant branch (AGB) phase, their presence is understood as a signature of mass accretion from an AGB companion that had since evolved into a white dwarf (WD; see, e.g., A. Jorissen et al., 2019, and references therein). During thermal pulses in the late-AGB phase, s-process elements such as barium and yttrium are mixed to the stellar surface and, if the AGB star fills its Roche Lobe, can be transferred to a binary companion. The detectability of these elements in the accretor’s photosphere depends on both the total amount of material accreted and the degree of dilution in the stellar envelope (L. Husti et al., 2009). Surface s-process enhancement, therefore, offers evidence of mass transfer, independent of the system’s current orbital configuration.
The third data release (DR3) of Gaia (Gaia Collaboration et al., 2023a) includes a catalog of non-single-star solutions (NSS; Gaia Collaboration et al., 2023b) for unresolved binaries, enabling astrometric orbit characterization across a wide range of orbital separations. S. Shahaf et al. (2024) identified MS+WD astrometric binaries in the NSS catalog, in which a faint WD accompanies a luminous MS star. These binaries have periods between days (corresponding to orbital separations of AU), and a substantial fraction of them exhibit moderate to high eccentricities ( up to 0.7). Standard binary population synthesis models do not predict this population, as they preferentially produce systems with shorter or longer orbital periods (see figure 16 of S. Shahaf et al. 2024 and discussion thereof). Its empirical properties and the processes that enabled it are the subject of ongoing studies (e.g., N. Hallakoun et al., 2024; N. Yamaguchi et al., 2024, 2025; O. Ironi et al., 2025; A. C. Rubio et al., 2025).
Identifying s-process enrichment in even part of this sample would link these systems to Barium stars, partially constrain the timing of mass transfer to the late- and post-AGB phase, and place them in a broader evolutionary context. To search for such signatures, P. Rekhi et al. (2024, R24 hereafter) combined this Gaia-selected sample with GALAH DR3 spectroscopy. They measured s-process abundances in 102 MS+WD binaries, and found that barium and yttrium enrichment in the MS primaries correlates with WD mass and metallicity. However, they did not find significant link with the orbital properties of the systems, possibly due to the limited sample size.
This study aims to extend the work of R24 by targeting eccentric MS+WD systems in the 100-1000 d period range. To achieve this, we conduct spectroscopic analysis to measure s-process abundances of 21 eccentric MS+WD binaries, complemented by a comparison sample of 12 systems with WD masses above 0.75 with nearly circular orbits. The targets were observed using the Fiber-fed Extended Range Optical Spectrograph (FEROS; A. Kaufer & L. Pasquini, 1998; A. Kaufer et al., 1999), mounted on the MPG/ESO 2.2-meter telescope at the La Silla Observatory in Chile. FEROS covers the spectral range from 350 to 920 nm at a resolving power of , making it well-suited for precise measurements of chemical abundances.
2 Sample Selection and Observations
2.1 Sample Selection
Our sample is drawn from the MS+WD candidates identified by S. Shahaf et al. (2024) using the astrometric triage method (see S. Shahaf et al., 2019, 2023). This method uses the astrometric mass-ratio function (AMRF) to link the photocentric orbit to the binary’s mass and flux ratios. Systems with AMRF values exceeding the maximum allowed by MS mass–luminosity relations cannot host a single MS companion and are classified as non-class-I binaries. The faint companions in these systems are likely WDs or close MS binaries. To reduce contamination from the latter, S. Shahaf et al. (2024) excluded systems with infrared excess. This process produced a sample of highly probable MS+WDs. The purity of this sample has been validated by several studies (e.g., N. Yamaguchi et al., 2024, 2025).
We selected two subsamples from the AMRF catalog for spectroscopic follow-up (Figure 1). The first consists of systems with WD masses of – and eccentricities . This mass range corresponds to the regime identified by R24 as favorable for s-process production during the AGB phase. Detecting enrichment in these systems would indicate past mass transfer when the progenitor radius was comparable to the current orbital separation, providing a test of whether high eccentricity can survive binary interaction or arise afterward. This region of parameter space was sparsely sampled in R24. This work extends it by targeting additional eccentric systems.
The second subsample is comprised from systems with WD more massive than and eccentricities below . These likely descend from progenitors with masses , which are expected to produce lower intrinsic s-process yields (A. I. Karakas & M. Lugaro, 2016). This high-M, low-eccentricity regime is absent from the R24 sample, and complementary in its properties to the first one.
Our targets were selected to be observable from La Silla, Chile, in January 2025, at airmass below 1.8 for sufficient time to reach predicted signal-to-noise ratios (SNRs) of per pixel at 4500 Å. We excluded systems already observed by GALAH and limited the sample to mag. These criteria respectively yielded 32 and 16 viable targets out of 529 and 41 total objects in the two categories described above, of which we selected the brightest 21 and 12 targets. 6 targets from the high-eccentricity subsample were also observed in an additional epoch in March 2025. Similar to the R24 sample, the systems in this work have near-solar primary masses, ranging between 0.8 to 1.2 .
2.2 Spectroscopic observations
Each target was observed with FEROS in consecutive exposures to assist in cosmic-ray removal; the exposures were co-added before analysis. The spectra were reduced using the ceres pipeline (R. Brahm et al., 2017), which provides continuum-normalized spectra. We masked normalized-flux values exceeding 1.2 or less than in order to avoid artifacts in the spectra. Due to particularly low SNR in our spectra below 4500 Å, this region was excluded from the analysis.
The target exposures were set to obtain an SNR per pixel of at 4500 Å based on the FEROS exposure time calculator111https://www.eso.org/observing/etc/bin/gen/form?INS.NAME=FEROS+INS.MODE=spectro (SNRs at longer wavelengths would be higher in FEROS spectra). However, the observed spectra had significantly lower SNRs around 4500 Å between 3.1 and 8.5, with a median SNR of 4.4. Nevertheless, most spectra had sufficient signal to identify the Fe and Ba lines required for our analysis. An example of a recorded spectrum with a typical SNR at the vicinity of the spectral lines analyzed in this work is shown in Figure 7.
3 Chemical abundances
The spectral modeling for fitting the metallicity and s-process element abundances was carried out using pymse (N. Piskunov & J. A. Valenti, 2017; A. Wehrhahn et al., 2023, v0.4.188). Spectral line data was obtained from the VALD3 database222https://vald.astro.uu.se/ (F. Kupka et al., 1999; N. E. Piskunov et al., 1995; T. Ryabchikova et al., 2015), with an atomic and molecular line list generated for a Sun-like star that included all lines with central depths . Using NLTE departure coefficient grids from A. M. Amarsi et al. (2020), pysme supports NLTE synthesis for Fe and Ba, however analysis for Y assumed LTE. To account for telluric features in the observed spectra, we employed telfit (K. Gullikson et al., 2014, v1.4.0), generating synthetic telluric models based on site-specific observational conditions, broadened to . Spectral regions where the modeled telluric transmission fell below 0.97 were masked from analysis.
We derived stellar parameters (, , [Fe/H]) and line-broadening velocities () from Fe lines selected for each star, following procedures adapted from S. Buder et al. (2021). Spectral fits were achieved by performing MCMC fits of pysme models to selected Fe lines, with local RV estimation and continuum detrending performed for every line segment. Seed values for , and [Fe/H] fits were taken from Gaia XP-based catalogs (R. Andrae et al., 2023; X. Zhang et al., 2023). A full description of the line selection, continuum treatment, RV corrections, and spectral fitting procedures is provided in Appendix A.
We identified 3 Ba lines (multiplets) — at 4554.03, 5853.68, 6496.90 Å in air — which are sufficiently deep and unblended with lines of other species across the range of stellar parameters of our sample. Similarly, we identified 3 Y lines — at 4900.12, 5087.42, 5200.41 Å in air. Ba and Y abundance values and their uncertainties were obtained through an MCMC analysis, with , , [Fe/H] and fixed to the previously estimated values. Local RV estimation and continuum detrending for each line segment followed the same procedure as for the Fe lines. An example of the spectral fit for Fe, Ba and Y abundances is shown in Figure 7. Due to the low SNR of the spectra, we excluded targets that did not have clearly distinguishable Ba and Y lines from the respective analysis. This excluded three targets from the Ba abundance analysis. As the Y lines are considerably shallower than the Ba lines at similar abundance levels, only 10 targets qualified for Y abundance analysis, all of them also having Ba abundance measurements.
Stellar parameters and abundances for the six two-epoch targets were derived by modeling both spectra in a joint MCMC framework, with each spectrum independently preprocessed (line selection, RV determination, continuum de-trending).
We obtained Fe and Ba abundances measured for 30 targets, 20 belonging to the (, M ) subsample, and 10 to the (M , ) subsample. Of these, 7 and 3 targets respectively have measured Y abundances. Following M. Van der Swaelmen et al. (2017) and R24, we define enrichment as cases where [Ba/Fe] or [Y/Fe] exceed 0.25 dex. A comparison of [Ba/Fe] and [Y/Fe] abundances is shown in Figure 9, confirming that Ba and Y enhancement typically occur together.
4 Results and Discussion
Figure 2 presents our measured Ba abundances as functions of eccentricity, WD mass and metallicity, with corresponding Y abundance plots shown in the appendix (Figure 10). As demonstrated in R24, as well previous Ba star studies (e.g., A. Jorissen et al., 2019; M. P. Roriz et al., 2021; B. Világos et al., 2024), the detection of s-process-enhancement depends strongly on metallicity and WD mass, which govern the efficiency of s-process nucleosynthesis in the WD progenitor (see M. Busso et al., 1999, 2001; A. I. Karakas & M. Lugaro, 2016). Our results align with this overarching paradigm while additionally highlighting the influence of eccentricity, as elaborated below.
4.1 AMRF MS+WD Systems as Ba-Stars
The results of R24 and this work show that, whenever metallicity permits, a substantial fraction of MS primaries in the AMRF sample exhibit Ba enhancement ([Ba/Fe] dex), consistent with expectations from AGB s-process yields.
Figure 3 compares the period-eccentricity distribution of the AMRF sample with literature samples of Ba dwarfs and giants (A. Escorza et al., 2019; A. Escorza & R. J. De Rosa, 2023) and analogous CH/CEMP-s stars (A. Jorissen et al., 2016). Both the overall AMRF distribution and the Ba-enhanced subset align with these populations, while extending coverage of Ba dwarfs to periods less than 1000 days (many with ), a region underrepresented in previous work. At these separations, the inflated radii of giant stars raise the possibility of dynamical interactions with their white dwarf companions. Although the current samples are too limited to draw statistical conclusions, the forthcoming Gaia DR4 release is expected to provide a larger sample of astrometric giant–white dwarf systems, enabling a more robust investigation of this effect.
Figure 4 further shows similar distribution of estimated WD masses (with MS companions) between the AMRF and the RV-based sample of A. Escorza & R. J. De Rosa (2023). The WD companions of Ba dwarfs have masses around 0.6 in both samples, indicating progenitor core development similar to field WDs. The generally higher WD masses in the Ba giants sample likely reflect a detection bias, as these systems tend towards higher primary masses.
These results indicate that the Gaia AMRF MS+WD binaries are products of interaction with AGB stars, likely through the same channels that produce Ba/CH/CEMP-s systems. Depending on metallicity, orbital period, and selection effects, they may appear as classical Ba stars. Gaia instead probes a complementary regime of shorter periods and a wider metallicity range, where enrichment is often weaker or harder to detect. Thus, even if not spectroscopically classified as s-process-enhanced, many of these systems are analogous. Notably, the current AMRF sample probes a volume more than an order of magnitude larger than the literature populations (Figure 5), providing a far broader view of the underlying parameter space. Future Gaia releases, extending to longer periods, will establish the AMRF sample as a benchmark for modeling binary interaction in intermediate-mass systems.
4.2 Eccentricity Dependence of Mass Transfer
In our sample of MS+WD binaries, both in this work as well as R24, we detect pronounced Ba and Y overabundance ([Ba/Fe], [Y/Fe] > 0.25) in many MS primaries up to the maximal orbital eccentricities probed (; Figures 2a and 10a). Such s-process enrichment clearly indicates past mass transfer—at least in the late and post-AGB phase—from the WD progenitor to the companion MS star; however, the persistence of significant eccentricity in these shows that tidal circularization, expected to accompany such mass transfer, either was much weaker than expected, or did not fully erase the orbital eccentricity.
However, this Ba enhancement seems to be sporadic, as only around half the systems with high eccentricity and sufficiently low metallicity demonstrate it. This points to as yet unconfirmed factors influencing mass-transfer in eccentric systems. One possibility is that these systems had different initial orbital parameters, perhaps being significantly more eccentric or more widely separated, reducing their accretion efficiency. Another possibility is that the progenitors of these WDs were much more massive than expected from the canonical single star initial-final mass ratio functions (J. D. Cummings et al., 2018; T. Cunningham et al., 2024), and had correspondingly lower s-process yields. Readers can refer to O. Ironi et al. (2025) and S. Shahaf (2025) for a detailed discussion.
We also identify Ba enhancement in four near-circular systems hosting relatively massive WDs (M ; Figure 2b). Three of these systems have [Fe/H] dex, a regime not probed in the R24 study. We suggest that the low metallicities of these systems mitigated the effect of their progenitors’ relatively high masses, allowing sufficient Ba production for their companions to accrete and exhibit enrichment. Given the straightforward explanation for the confirmed s-process enhanced high-M systems, it is difficult to assess whether their (current) low eccentricity corresponds to higher mass transfer efficiency. Nevertheless, our sample extends the empirical evidence for mass transfer in these systems to encompass higher WD masses.
4.3 Formation Channels
Our results add to a growing body of work documenting significant eccentricities in various classes of post-interaction binaries (e.g., A. I. Karakas et al., 2000; A. Jorissen et al., 2016, 2019; J. Vos et al., 2017; M. Van der Swaelmen et al., 2017; G.-M. Oomen et al., 2018; J. Müller-Horn et al., 2025). A particularly relevant parallel is provided by G.-M. Oomen et al. (2018), who studied a sample of 33 post-AGB+MS binaries with orbital periods of days, most of which exhibit significant eccentricity (). Our work demonstrates that these eccentricities persist through the post-AGB phase to the WD formation, diminishing the possibility that these systems could yet circularize via accretion and/or interaction with the post-AGB planetary nebula or circumbinary (CB) disk. Tidal evolution theory predicts that binaries in which one component fills its Roche lobe should circularize on relatively short timescales, far shorter than the expected duration of the mass transfer phase in the progenitor system (see, e.g., S. F. Portegies Zwart & F. Verbunt, 1996; G. Nelemans et al., 2001; J. F. Sepinsky et al., 2007, 2009; S. Toonen et al., 2012; S. Toonen & G. Nelemans, 2013). Multiple models have been proposed to explain this seemingly anomalous population of eccentric post-interaction systems. Broadly, these fall into two categories: (1) Mechanisms that avoid strong tidal circularization altogether; (2) Mechanisms that actively pump eccentricity, counteracting canonical tidal damping.
In the first category, one possibility is that tidal forces are much weaker than expected in certain systems (e.g., J. D. Nie et al., 2017), allowing the Roche-lobe overflow (RLOF) phase to conclude before circularization can occur. This seems unlikely for our sample, except perhaps in extremely eccentric cases, because in most systems at these separations, the WD progenitor would have substantially filled its Roche lobe for much of the RGB and the entirety of the AGB phase, thereby experiencing significant tidal forces (J. P. Zahn, 1977). Another possibility is that mass transfer proceeds primarily via stellar winds, rather than RLOF, which can produce non-trivial eccentricity evolution (H. M. J. Boffin & A. Jorissen, 1988; H. M. J. Boffin, 2015; M. I. Saladino et al., 2018; M. I. Saladino & O. R. Pols, 2019; S. L. Schrøder et al., 2021). A third option is that the observed eccentricity reflects the presence of a tertiary companion, either through direct orbital perturbations or because the tertiary itself was the AGB donor, later being ejected or migrated to a wider orbit after the interaction (H. B. Perets & K. M. Kratter, 2012).
To investigate the third possibility, we employed the algorithm of K. El-Badry et al. (2021) to search for wide companions to our systems using their NSS-derived astrometric parameters. No tertiary matches were found for any of the s-process-enriched systems reported here or in R24. However, the detection probability declines sharply for companions fainter than , which, for our sample, effectively excludes all WDs with cooling ages older than a Gyr as well as the majority of M dwarfs. A more definitive assessment of this scenario will thus require deeper surveys, such as future Gaia data releases and surveys by the Vera Rubin Observatory (Ž. Ivezić et al., 2019).
In the second category, a frequently invoked eccentricity-pumping mechanism is the formation of a CB disk from AGB ejecta, with resonant disk–binary interactions driving eccentricity growth (T. Dermine et al., 2013; J. Vos et al., 2015; D. J. D’Orazio & P. C. Duffell, 2021; M. Siwek et al., 2023; R. Valli et al., 2024). Other proposed mechanisms include phase-dependent mass loss via RLOF and/or winds (F. Dosopoulou & V. Kalogera, 2016; A. Parkosidis et al., 2025), and tidally enhanced mass loss at periastron (A. A. Bonačić Marinović et al., 2008; A. Kashi & N. Soker, 2018). As demonstrated by A. Parkosidis et al. (2025), phase-dependent RLOF mass-loss can cause eccentricity-pumping, but comprehensive population synthesis modeling is required to determine whether they can reliably reproduce the orbital characteristics presented here and in other studies.
However, we note that many existing studies often yield contradictory results or fail to explore the full parameter space relevant to our systems, and those that especially explore the d period range covered by our sample have been unable to reproduce the observed eccentricities (R. R. Rafikov, 2016; M. I. Saladino & O. R. Pols, 2019; G.-M. Oomen et al., 2020; P. Krynski et al., 2025). The recent work by P. Krynski et al. (2025) presented detailed models of the evolution of AGB+MS binaries, incorporating many of the mechanisms discussed above, including classical and wind RLOF, tidally enhanced winds, non-conservative mass transfer, and eccentricity pumping due to a CB disk. Even under favorable assumptions, however, the models failed to reproduce the observed eccentricities of post-interaction MS+WD binaries with periods shorter than days. This discrepancy indicates that key aspects of the underlying processes remain poorly understood.
5 Outlook
The clear detection of Ba-enhanced MS+WD systems despite the relatively low SNR of our spectra underscores the strength of the signal and motivates deeper follow-up observations. Higher-SNR () spectra, even at lower resolving power, would enable precise abundance determinations for individual systems, and hence constrain the efficiency of mass transfer as a function of orbital parameters, WD mass and metallicity.
We propose a survey of MS+WD systems from the catalog of S. Shahaf et al. (2024) using the HighSpec spectrograph (Y. Sofer Rimalt et al., 2024) on the upcoming Multi-Aperture Spectroscopic Telescope (MAST; Ben-Ami et al., in prep) array at Neot Smadar, Israel, with a particular focus on systems exhibiting sub-solar metallicities and moderate eccentricities. HighSpec employs a custom narrow bandpass ion-etched grating at and achieves an end-to-end efficiency of 55%, making it well suited for observing prominent spectral lines even in relatively faint targets. Our primary objective is to measure the strong Ba II absorption line at 4555 Å. The surrounding Å band also includes numerous relatively strong Fe lines, enabling precise determination of [Fe/H]. Figure 6 shows the limiting magnitude achievable at various SNRs as a function of exposure time, overlaid on the JKC -band cumulative magnitude distribution of MS+WD systems with sub-solar metallicities (as per R. Andrae et al. 2023) accessible from Israel. With up to 20 MAST telescopes operating simultaneously with HighSpec, we can observe of the brightest targets (including roughly half with ) in nights.
Appendix A Determination of RV, , , [Fe/H] and
The determination of effective temperature (, surface gravity (), iron abundance ([Fe/H]) and rotational broadening velocity was carried out using a set of Fe lines custom-selected for each star. As the low SNR of our spectra makes simultaneous fitting of projected rotational velocity () and macro-turbulence velocity () challenging, we followed S. Buder et al. (2021) and fit for (defined as with set to 0) instead. The microturbulence velocity () was estimated using the empirical relation provided in the GALAH DR3 study (eqns 4 and 5; S. Buder et al., 2021). During the following procedure, where required, initial/seed values of the effective temperature (), surface gravity () and iron abundance ([Fe/H]; here used synonymously with metallicity) for each star were derived from Gaia’s low-resolution XP spectra using the XGBoost algorithm (R. Andrae et al., 2023).
For each selected line, we extracted 1 Å-wide spectral segments centered on the line, merging overlapping segments. Lines were chosen based on the following criteria: (1) the SNR in the line region exceeded 10, except for three stars where the threshold was lowered to 5 due to fewer than five lines meeting the SNR condition; (2) the Fe line depth would exceed 0.4 at [Fe/H] = 0 for the star’s and ; (3) there were no lines of other species within Å exceeding 20% of the Fe line’s depth; and (4) the mean value of the continuum trend (see next) across the Å segment was greater than 0.9, with a maximum variation of less than 0.02. The fourth condition excluded regions where the blaze function caused the pipeline’s continuum normalization quality, as well as the local SNR, to drop considerably. To reduce computational expenses, when the number of such segments exceeded 30 (for 5 spectra), we selected the 30 segments with the deepest lines.
To correct for the local continuum trend, we fit a smoothing spline to continuum points in 60 Å windows centered on each segment of interest using the scipy function scipy.interpolate.make_smoothing_spline. For this procedure, synthetic spectra were created with stellar parameters as described above and Doppler shifted to stellar rest frames using the nominal radial velocity (RV) determined using the H line. Continuum points were denoted as those where the model flux exceeded 0.98, which ensured a sufficient number of points for robust spline fits with negligible bias from absorption features.
The relative uncertainty in the spectrum in each segment was estimated from the standard deviation of the continuum-identified points in a 20 Å window centered on the segment. To isolate the noise estimate from the continuum trend, the normalized flux in the windows was first divided by the continuum spline.
To determine the local RV value around each line segment, we used a 20 Å-wide region centered on the corresponding 1 Å segment. Regions (and corresponding 1 Å segments) with more than 25% of the flux points masked were discarded to ensure reliability both in RV estimation and continuum de-trending. Synthetic spectra were generated using the given stellar parameters, including only lines with absorption depths greater than 0.1. The RV was determined using cross-correlation with a built-in method from pymse. A clipping procedure was then used to identify and discard any windows with unreliable RVs, typically caused by an insufficient number of deep lines in the region. Note that these estimated RV values are not corrected for Earth’s orbital motion, as this is not required for our science case. Across all spectra, the mean (max) range in the measured RVs across all usable windows is km s-1, and the mean (max) standard deviation is km s-1.
Each 1 Å segment was then Doppler-shifted to Earth’s rest frame using its locally determined RV value, rather than a global mean, to avoid potential inter- and intra-order wavelength calibration drifts. Subsequently, segments with any masked data points within the core line region — defined as a Å interval around the line center — were removed from further analysis. After this final filtering step, between 6 and 30 segments per spectrum remained, with a mean of 21 segments per spectrum.
Finally, we conducted a Markov Chain Monte Carlo (MCMC) analysis to fit for , , [Fe/H] and , treating all segments simultaneously. For this we created a bank of synthetic spectra using pysme on a grid of , , [Fe/H] values, at steps of 100 K, 0.1 dex and 0.05 dex respectively, with a wavelength resolution of 0.003 Å. The line list passed to pysme during the synthesis of this bank was restricted to only Fe lines in the relevant segments in order to avoid prohibitively large computational times, as computation time scaled linearly with the number of lines in the list. This restriction does not affect our results, as the Fe lines were previously selected to avoid significant blending with lines from other species. For each star, we adopted weakly informative normal priors for and , centered on their seed values with conservative standard deviations of 250 K and 0.25 dex respectively, and uniform priors on [Fe/H] and spanning dex and km s-1 respectively. During each MCMC iteration, unbroadened synthetic spectra at the given three stellar parameters were interpolated from the grid and were rotationally and instrumentally broadened using the routines fastRotBroad, with limb-darkening parameter set at 0.81, and instrBroadGaussFast from pyastronomy (version 0.22.0; S. Czesla, 2025). Instrumental resolution at the center of each segment was computed from the wavelength solution, assuming FEROS’s sampling of 2.2 pixels per resolution element (A. Kaufer & L. Pasquini, 1998). We compared spectra broadened with pysme and pyastronomy and verified that the maximum deviation between the two was less than 1%. The broadened spectra were then interpolated onto the FEROS wavelength grid and multiplied by the continuum trend. We used as the log-likelihood function, calculated from all data points from each Fe segment selected for that star. An example of the spectral fit in Figure 7a.
Appendix B Additional Plots
Figure 8 shows Ba abundance as a function of [Fe/H] and M as a complementary plot to figure 3 of R24. Y abundance plots that complement Figure 2 are presented in Figures 9 10.
Appendix C Data Tables
Gaia DR3 ID | Period (d) | Eccentricity | M () | M1 () | [Fe/H] | [Ba/Fe] | [Y/Fe] |
3036788173176962048 | – | ||||||
4790658375100436736 | |||||||
3208829887146528384 | |||||||
6091100873778091392 | |||||||
5620705644518714752 | |||||||
5406806102703951360 | |||||||
3772123820203876736 | – | ||||||
5397347794427206656 | |||||||
6181668466886465280 | |||||||
2914005812780342272 | – | ||||||
5722712938857848192 | – | ||||||
5164360455867611264 | – | ||||||
2923820637686349568 | – | ||||||
4795585179328518016 | |||||||
5550946678313327744 | – | ||||||
6169594184243050752 | – | ||||||
5327240562221030528 | – | ||||||
2964151136511557120 | – | ||||||
5317828124219671168 | – | ||||||
6300835156308377344 | |||||||
5329433194583181952 | – | ||||||
2919286664051008640 | – | ||||||
6192821814055372800 | – | ||||||
5744310130364616576 | – | ||||||
3697174751104247808 | – | ||||||
5890631694717361408 | – | ||||||
3147042801161881472 | – | ||||||
2923409729576118912 | – | ||||||
5091887624392418816 | – | ||||||
3633619813626022784 |
Gaia DR3 ID | Period (d) | Eccentricity | M () | M1 () | [Fe/H] | [Ba/Fe] | [Y/Fe] |
6472197475176313472 | |||||||
6645994082722928768 | |||||||
4878036323641744384 | |||||||
4646881493008180352 | |||||||
4774266821393569792 | |||||||
6489348928855798272 | |||||||
5495572631335262208 | |||||||
5707071630039205248 | |||||||
4715941990147018752 | |||||||
4881348087024892416 | |||||||
5363208336349003648 | |||||||
2436150255390344320 | |||||||
3623765616756619392 | |||||||
6669412928776138112 | |||||||
4693571910684108928 | |||||||
6081919436288801024 | |||||||
5435770125183875200 | |||||||
2600643791975931264 | |||||||
2545472433093014912 | |||||||
4650039354829199616 | |||||||
3285029929897944320 | |||||||
4622751301587564416 | |||||||
5371201888966580992 | |||||||
5386048560107201792 | |||||||
6402712399045698560 |